# MoTCoder: Elevating Large Language Models with Modular of Thought for Challenging Programming Tasks

### 26 Dec 2023 v1

@@ 3.1 Modular-of-Thought Instruction Evolution
The instruction is supplemented with a one-shot example, serving to prompt the model to adhere to the MoT instruction generation strategy. An illustration of the instruction prompt is presented in Figure 3 and further detailed in the Appendix. The instruction generated by MoTCoder encourages the model to decompose a program into sub-modules. This mirrors the methodology commonly employed by developers when addressing intricate coding tasks, where they systematically break down solutions into modular components.
@@
(a) Modular Prompt
(b) Normal Prompt
Evolution Examination.
  1. 1.
  2. 2.
  3. 3.
3.2 Modular-of-Thought Training and Inference
Table 4: APPS test results by pass@k (%). Competitive approaches include the fine-tuned GPT-Neo (Hendrycks et al., 2021a), one-shot Codex (Chen et al., 2021c), fine-tuned AlphaCode (Li et al., 2022c), one-shot StarCoder (Li et al., 2023), one-shot WizardCoder (Luo et al., 2023b) and two-shot CodeLlama series models (Roziere et al., 2023).
@@ 3.2 Modular-of-Thought Training and Inference
GPT-Neo 2.7B 1 3.90 0.57 0.00 1.12
5 5.50 0.80 0.00 1.58
Codex 12B 1 4.14 0.14 0.02 0.92
5 9.65 0.51 0.09 2.25
1000 25.02 3.70 3.23 7.87
AlphaCode 1B 1000 17.67 5.24 7.06 8.09
AlphaCode (Filtered 1k) 5 14.36 5.63 4.58 7.17
AlphaCode (Filtered 10k) 5 18.18 8.21 6.65 9.89
AlphaCode (Filtered 50k) 5 20.36 9.66 7.75 11.42
@@ 4.3.1 Results on APPS
We conducted a comparison of our approach with existing LLM baselines on APPs (Hendrycks et al., 2021a). All outcomes are computed using raw predictions without being filtered by the test cases provided in the prompt. We juxtapose our findings with GPT-Neo (Hendrycks et al., 2021a), Codex (Chen et al., 2021c), AlphaCode (Li et al., 2022c), StarCoder (Li et al., 2023), WizardCoder (Luo et al., 2023b) and CodeLlama series models (Roziere et al., 2023).
As depicted in the results from the APPS shown in Table 4, we notice that the modular-of-thought training approach leads to an enhancement in code generation capabilities comparing with previous instruction finetuning models. Our MoTCoder exhibits improved performance across all difficulty levels and demonstrates more substantial gains in the interview and competition-level problems, characterized by more intricate solutions. To provide specifics, the pass@1 performance of MoTCoder-15B surpasses the pass@5 performance of CodeLlama-Python-34B, showcasing significant improvements of 11.9% on interview-level problems and 7.6% on competition-level problems, despite fewer model parameters and inference times. Moreover, our pass@1 on competition-level questions surprisingly outperformed the closed-source model GPT3.5 by an impressive margin of 5.7%. This outcome serves as compelling evidence of the efficacy of our approach in addressing intricate problems.

### 05 Jan 2024 v2

@@ 1 Introduction
Figure 4: Illustration of modular code generation involves the model initially delineating the necessary sub-modules, each comprising a function header and a docstring detailing its intended purpose. Following this, the model proceeds to comprehensively implement each module in code and incorporates them as integral components of the finalized solution..
@@ 3.1 Modular-of-Thought Instruction Evolution
The instruction is supplemented with a one-shot example, serving to prompt the model to adhere to the MoT instruction generation strategy. An illustration of the instruction prompt is presented in Figure 3 and further detailed in Figure 4. The instruction generated by MoTCoder encourages the model to decompose a program into sub-modules. This mirrors the methodology commonly employed by developers when addressing intricate coding tasks, where they systematically break down solutions into modular components.
(a) Modular Prompt
(b) Normal Prompt
Evolution Examination.
  1. 1.
  2. 2.
  3. 3.
3.2 Modular-of-Thought Training and Inference
@@
Table 4: APPS test results by pass@k (%). Competitive open-source approaches include CodeT5 (Wang et al., 2021), fine-tuned GPT-Neo (Hendrycks et al., 2021a), GPT-2 (Radford et al., 2019), GPT-3 (Brown et al., 2020), one-shot StarCoder (Li et al., 2023), and one-shot WizardCoder (Luo et al., 2023b). We also present results from closed-source models, namely text-davinci, code-davinci, and GPT3.5.
@@
CodeT5 770M 1 6.60 1.03 0.30 2.00
GPT-Neo 2.7B 1 14.68 9.85 6.54 10.15
5 19.89 13.19 9.90 13.87
GPT-2 0.1B 1 5.64 6.93 4.37 6.16
5 13.81 10.97 7.03 10.75
1.5B 1 7.40 9.11 5.05 7.96
5 16.86 13.84 9.01 13.48
GPT-3 175B 1 0.57 0.65 0.21 0.55
@@ 4.3.1 Results on APPS
We conducted a comparison of our approach with existing large language model baselines on APPs (Hendrycks et al., 2021a). All outcomes are computed using raw predictions without being filtered by the test cases provided in the prompt. Our analysis includes a comparison with open-sourced approaches such as CodeT5 (Wang et al., 2021), fine-tuned GPT-Neo (Hendrycks et al., 2021a), GPT-2 (Radford et al., 2019), GPT-3 (Brown et al., 2020), one-shot StarCoder (Li et al., 2023), and one-shot WizardCoder (Luo et al., 2023b). Additionally, we present results from closed-source models, including text-davinci, code-davinci, and GPT3.5.
As depicted in the results from the APPS shown in Table 4, we notice that the modular-of-thought training approach leads to an enhancement in code generation capabilities compared with previous instruction finetuning models. Our MoTCoder exhibits improved performance across all difficulty levels and demonstrates more substantial gains in the interview and competition-level problems, characterized by more intricate solutions. To provide specifics, the pass@1 performance of MoTCoder-15B surpasses WizardCoder-15b by 15.49% on interview-level problems and 10.28% on competition-level problems. Moreover, our pass@1 on competition-level questions surprisingly outperformed the closed-source model GPT3.5 by an impressive margin of 5.7%. This outcome serves as compelling evidence of the efficacy of our approach in addressing intricate problems.